TENTATIVE Program for the 1st International IMACS Conference on Applications of Computer Algebra

May 16, 1995 (Tuesday)

Session Symmetries and Closed Form Solutions of Differential Equations Fritz Schwarz SUB 250-A Session Parametric Curves and Applications in Computer Aided Geometric Design J. Rafael Sendra

SUB 231-A

8:30 am Group-Invariant Solutions of Hydrodynamics Stephen Coggeshall Parametric Curves and Applications in CAGD: Introduction and Some Special Issues Tomas Recio, J. Rafael Sendra* & Juana Sendra

9:00 am Symmetry Reduction and Symbolic Solutions of Ordinary Differential Equations Alexei Bocharov Implicitization by Groebner Basis Conversion Michael Kalkbrener

9:30 am JET—An Axiom Environment for Geometric Computations with Differential Equations Werner Seiler & Jacques Calmet* New Elimination Tools based on Symmetric Functions for the Implicitization of Parametric Curves and Surfaces Laureano Gonzalez-Vega

10:00 am  BREAK

10:30 am The Characterization of 3rd Order ODE’s Admitting Fiber Preserving Point Symmetries Guy Grebot Issues in Symbolic Parametrization of Algebraic Curves Franz Winkler

11:00 am To Be Announced Thomas Wolf Geometric Constraint Solving: (I) Graph-Directed Solvers Christoph M. Hoffmann

11:30 am Symmetries and Solutions of Plasma-Fluid Equations Raul Acevedo A Toolkit for Algebraic Geometry Ashutosh Rege

12:00 pm  LUNCH

Session Symmetries and Closed Form Solutions of Differential Equations Fritz Schwarz SUB 250-A Session Polynomial Systems

Lakshman Y. N. & David Wood SUB 231-A

1:30 pm Symmetries for Dynamical Systems on a Lattice Decio Levi Geometric Constraint Solving: (II) Algebraic Issues Christoph M. Hoffmann

2:00 pm Resurrecting Dixon Resultants

Deepak Kapur* & Tushar Saxena

2:30 pm Algorithmic Determination of the Cartan Structure of Infinite Lie Symmetry Groups of PDE’s Greg Reid The Kapur-Saxena-Yang Dixon Resultant with Maple and Mathematica Janet McShane, George Nakos & Robert M. Williams

3:00 pm  BREAK

3:30 pm Symbolic Computation and Conserved Densities Willy Hereman On Sparsifying Transformations for Multivariate Polynomials Lakshman Y. N.

4:00 pm Liouvillian Solutions of 3rd Order ODE’s: A Progress Report Winfried Fakler & Jacques Calmet* Quasi-Linear Differential Algebraic Equations: An Algebraic Study Gabriel Thomas

4:30 pm On the Translation Invariants of Quadratic Differential Systems Leonid A. Timochouk Geometry and Polynomial Systems—Theory and Application Peter F. Stiller

5:00 pm Ten Minutes of Discussion of Issues Raised and Not Raised informally led by the organizers

evening  Sacrifice at the flying saucer landing pad

May 17, 1995 (Wednesday)

Session Construction and Applications of Normal Forms in the Theory of Nonlinear ODEs and Transfer Map Techniques Alexander D. Bruno & Victor Edneral SUB 250-A Session Computer Algebra Applied to Technology & Open Session

Waldir L. Roque & Laureano Gonzalez-Vega SUB 231-A

8:30 am Methods of calculation of normal forms Alexander D. Bruno Mixing Symbolic and Numeric Strategies to Solve the Inverse Kinematics Problem in Robotics Laureano Gonzalez-Vega

9:00 am A Lie Algebraic Approach Jan Sanders Design and Performance Analysis of Optical Instruments with Computer Algebra Yves Papegay

9:30 am Normal Form Method and Formal Integrals of Nonlinear ODEs systems Victor F. Edneral Verification of Knowledge Based Systems with Commutative Algebra and Computer Algebra Techniques Eugenio Roanes-Lozano* & Luis M. Laita

10:00 am  BREAK

10:30 am SAP, a symbolic algebraic processor Andre Deprit Two more links to NAG numerics involving CA systems Evatt Hawkes & Grant Keady*

11:00 am Computation of coefficients of the normal form in a singular case Sergey Yu. Sadov A Review of CAS Mathematical Capabilities Michael Wester

11:30 am Phase space structure of 4D symplectic mappings and applications to beam optics Ezio Todesco On Conservation Laws and Symmetries for PDEs Vladimir Rosenhaus

12:00 pm  LUNCH

Session Polynomial Systems Lakshman Y. N. & David Wood SUB 250-A Session CA Applications of Stochastic Methods

Michael Trott

SUB 231-A

1:30 pm Matrix Methods for Sparse Polynomial Systems John Canny Quantum coupling coefficients seen as discrete wave functions Markus van Almsick

2:00 pm Advances in Affine Elimination Theory J. Maurice Rojas Computer-Aided Stochastic Calculus with Applications in Finance Colin P. Williams

2:30 pm Polynomial Conditions on Spectra of Real, Symmetric Toeplitz Matrices David H. Wood Algebra for the Electronic Marketplace Ross Miller

3:00 pm  BREAK

3:30 pm A Normal Form Algorithm for Matrices over k[x,y]/<xy> Reinhard Laubenbacher Transmission of Solitons through Random Media Robert Knapp

4:00 pm Ten Minutes of Discussion of Issues Raised and Not Raised informally led by the organizers

evening  Tramway ride to Sandia Peak & dinner

May 18, 1995 (Thursday)

Session Parallel Computer Algebra Eric Kaltofen & Gilles Villard Woodward Hall 147 Session Numerical Methods for PDEs in a Computer Algebra Environment Victor G. Ganzha & E. V. Vorozhtsov

Woodward Hall 149

8:30 am REDUCE on a Massively Parallel System Winfried Neun Applications of Computer Algebra in Scientific Computing Hans-Joachim Bungartz

9:00 am Parallel Computer Algebra on the Desk-Top Wolfgang Kuechlin & Beatrice Amrhein* Symbolic-Numeric Stability Investigation of Difference Schemes for the Euler and Navier-Stokes Equations Victor G. Ganzha & E. V. Vorozhtsov

9:30 am Parallelism in MuPAD Holger Naundorf Using the Symbolic Computation for Analysis of Local Approximation of Finite-Difference Operators Mikhail Shashkov* & Victor G. Ganzha

10:00 am  BREAK

10:30 am A Flexible Tool to Experiment with Distributed Symbolic Computation Stephane Dalmas & Marc Gaetano* Algorithm for Solving Riccati Matrix Equation with Guaranteed Accuracy Aider Ya. Bulgakov

11:00 am The Newton Polyhedra in the Nonlinear Analysis Alexander D. Bruno Local Approximation Study of Johnson’s Method for 2D Elasticity Problems Victor G. Ganzha & E. V. Vorozhtsov

11:30 am FORUM

started off by the conference organizers

12:00 pm  LUNCH

Session Parallel Computer Algebra Eric Kaltofen & Gilles Villard Woodward Hall 147 Session CA Methods in Control

Chaouki T. Abdallah & Peter Dorato

Woodward Hall 149

1:30 pm Parallel Integral Multivariate Polynomial Greatest Common Divisor Computation Mohamed Omar Rayes Complexity of Systems and Control Theory Problems Paul W. Goldberg, Chaouki T. Abdallah* & Greg L. Heileman

2:00 pm Distributed Symbolic Computation with the Black Box Representation Angel Diaz* & Erich Kaltofen NP-hardness of some linear control design problems Vincent Blondel* & John N. Tsitsiklis

2:30 pm Parallel Quantifier Elimination on the SP2 Hoon Hong*, Michael Jahn, Richard Liska, Nicholas Robidoux & Stanly Steinberg Application of Quantifier Elimination Theory to Robust Multi-objective Feedback Design Peter Dorato*, Wei Yang & Chaouki T. Abdallah

3:00 pm Summarizing Remarks

Erich Kaltofen

3:30 pm  BREAK

4:00 pm BUSINESS MEETING

all who wish to attend

evening  Social gathering at Professor Steinberg’s house

May 19, 1995 (Friday)

Session CAS in Engineering Education H. Eric Nuttall & Jeanine Ingber SUB 250-A Session Computer Algebra and Automated Theorem Proving

Dongming Wang

SUB 231-A

8:30 am An Open Environment for Doing Mathematics

Karsten Homann & Jacques Calmet

9:00 am Geometry Theorem-Proving and Geometry Problem-Solving, Old and New Wen-tsun Wu

9:30 am Radical Decomposition and Geometry Theorem Proving

Michael Kalkbrener

10:00 am  BREAK

10:30 am Coding Theory in MATHEMATICA Igor Gachkov & Kenneth Hulth Zero Decomposition for Theorem Proving in Geometry Dongming Wang

11:00 am Geometric Reasoning and Resultants

Deepak Kapur

11:30 am Probabilistic Verification of Elementary Geometry Statements Giuseppa Carra Ferro, Giovanni Gallo* & Rosario Gennaro

12:00 pm  LUNCH

Session CAS in Engineering Education H. Eric Nuttall & Jeanine Ingber SUB 250-A Session Quantifier Elimination and its Applications

Richard Liska & Michael Jahn

SUB 231-A

1:30 pm Advanced Engineering Math Through Symbolic Software Robert Lopez To Be Announced Hoon Hong

2:00 pm Active learning through interactive texts based on Mathematica Notebooks Jerry Uhl Partial solution of a path finding problem using the CAD method Scott McCallum

2:30 pm Engineering Applications of Computer Algebra: Mathematica Approach Cetin Cetinkaya Stability Analysis by Quantifier Elimination Richard Liska* & Stanly Steinberg

3:00 pm  BREAK

3:30 pm Freshmen Engineering: Computing using Mathematica Jeanine Ingber & Eric Nuttall Using Quantifier Elimination to solve the Birkhoff Interpolation Problem Laureano Gonzalez-Vega

4:00 pm Mathematica for the Beginning Teacher Henry Shapiro Applying quantifier elimination to problems in optimization and simulation Volker Weispfenning